Supply the missing data in the following cases. Each case is independent of the others. (Leave no cells blank - be certain to enter O wherever required.) Case 2 4 Direct materials $ 4,700$6,200S 5,200$ 3,200 4,200 9,200 Direct labour 3,200 7,200 5,200 18,700 2,700 Manufacturing overhead Total manufacturing costs Beginning work-in-process invento Ending work-in-process inventory Cost of goods manufactured Sales 4,200 20,200 3,200 4,200 1,200 14,500 22,000 2,700 3,100 18,200 30,000 1,200 40,200 2,200 17,900 36,200 Beginning finished goods inventory Cost of goods manufactured Goods available for sale 4,200 17,000 19,200 Ending finished goods inventory 1,700 3,700 Cost of goods sold Gross margin Operating expenses Net income 16,000 14,000 3,650 $4,200 $ 5,4009,200
Answer #1

Case

1

2

3

4

Direct Material

$    4,700.00

$    6,200.00

$    5,200.00

$    3,200.00

Direct labor

$    8,800.00

$    3,200.00

$    7,200.00

$    4,200.00

Manufacturing overheads

$    5,200.00

$    4,200.00

$    7,800.00

$    9,200.00

Total Manufacturing cost

$ 18,700.00

$ 13,600.00

$ 20,200.00

$ 16,600.00

Beginning WIP

$    2,700.00

$    2,100.00

$    3,200.00

$    4,400.00

Ending WIP

$    3,200.00

$    1,200.00

$    4,200.00

$    3,100.00

Cost of goods manufactured

$ 18,200.00

$ 14,500.00

$ 19,200.00

$ 17,900.00

Sales

$ 30,000.00

$ 22,000.00

$ 36,200.00

$ 40,200.00

Beginning finished goods inventory

$    1,200.00

$    2,700.00

$    2,000.00

$    2,200.00

Cost of goods manufactured

$ 18,200.00

$ 14,500.00

$ 19,200.00

$ 17,900.00

Goods available for sale

$ 19,400.00

$ 17,200.00

$ 21,200.00

$ 20,100.00

Ending finished goods inventory

$    3,400.00

$    1,700.00

$    4,200.00

$    3,700.00

Cost of goods sold

$ 16,000.00

$ 15,500.00

$ 17,000.00

$ 16,400.00

Gross profit

$ 14,000.00

$    6,500.00

$ 19,200.00

$ 23,800.00

Operating expenses

$    9,800.00

$    3,650.00

$ 13,800.00

$ 14,600.00

Net income

$    4,200.00

$    2,850.00

$    5,400.00

$    9,200.00

Working

Case

1

2

3

4

Direct Material

4700

6200

5200

3200

Direct labor

=18700-4700-5200

3200

7200

4200

Manufacturing overheads

5200

4200

=20200-7200-5200

9200

Total Manufacturing cost

18700

=6200+3200+4200

20200

=3200+4200+9200

Beginning WIP

2700

=14500+1200-13600

3200

=17900+3100-16600

Ending WIP

=18700+2700-18200

1200

4200

3100

Cost of goods manufactured

18200

14500

=20200+3200-4200

17900

Sales

30000

22000

36200

40200

Beginning finished goods inventory

1200

2700

=21200-19200

2200

Cost of goods manufactured

18200

14500

19200

17900

Goods available for sale

=1200+18200

=2700+14500

=17000+4200

=2200+17900

Ending finished goods inventory

=19400-16000

1700

4200

3700

Cost of goods sold

16000

=17200-1700

17000

=20100-3700

Gross profit

14000

=22000-15500

19200

=40200-16400

Operating expenses

=14000-4200

3650

=19200-5400

=23800-9200

Net income

4200

=6500-3650

5400

9200